d = ℓ\sqrt{N}

τ = ρκs

\frac{dI}{I} = −ρκ ds = −dτ

I(τ) = I_{0} e^{−τ}

dI = −ρκ ds I

−\frac{1}{κρ} \frac{dI}{ds} = I

dI = −ρκ ds I + ρj ds

\frac{1}{κρ} \frac{dI}{ds} = \frac{j}{κ} − I = S − I

I(s) = ∫_{s=0}^{s=s_{f}} \frac{1}{κρ} \frac{dI}{ds} = ∫ \frac{j}{κ} − I = S − I

I(s) = I_{0} e^{−ρκs} + S (1 − e^{−ρκs})

\frac{I(θ)}{I(θ = 0)} = a + b cos(θ)